OpenStudy (anonymous):

write an algebraic expression that is equivalent to the expression. sec[arcsin(x-1)]

myininaya (myininaya):

\[\sec(\arcsin(x-1))\] let y=arcsin(x-1) => sin(y)=x-1 sin(y)=(x-1)/1 sin(y)=opp/hyp 

myininaya (myininaya):

Find the other side in terms of x using the Pythagorean thm

myininaya (myininaya):

then since y=arcsin(x-1) then using the triangle above to answer sec(y) will be your answer

myininaya (myininaya):

sec(y)=hyp/adj remember? you already have hyp=1 you need the adj side which can be found by the Pythagorean thm like I said above

OpenStudy (anonymous):

Thanks!!